# Force-free wave interaction in magnetar magnetospheres: Computational modeling in axisymmetry

### v1

@@ 4 Discussion
Here, $\mu_{s}\approx 10^{30}\,{\rm erg}\,{\rm cm}^{-3}$ is the shear modulus of the deep crust and $s<0.1$ denotes the elastic strain. AWs emerging from the stressed regions during the quake duration $t_{\rm Quake}$ carry a fraction $f$ of $E_{\rm Quake}$ (Bransgrove et al., 2020). Their power is:
@@
Throughout Section 3 we numerically evaluate conversion efficiencies $\eta$ for two different mode conversion mechanisms in axisymmetry: mode conversion at curved magnetic field lines and by non-linear three-wave interactions. Typical conversion efficiencies range between $\eta\approx 1-10\%$ for low-amplitude waves; higher efficiencies can be achieved by counter-propagating interacting AWs and for large seed wave amplitudes (see Section 3.3). Directly applying these efficiencies to Equation (25) and assuming that most of the FMS wave energy is concentrated in the short initial pulse of the first magnetospheric crossing yields FMS wave luminosities of $L_{\rm FMS}\approx 10^{42}-10^{43}\,{\rm erg}/{\rm s}$. AWs can generate FMS waves of twice their frequency during the first crossing of the magnetosphere. Subsequent bounces produce more FMS waves during each crossing. These late-episode FMS waves gradually decrease in luminosity but extend the possible signal durations substantially. We describe the FMS wave tail generated by ongoing AW activity in the inner magnetosphere by a fit function (see Appendix A). Especially for the long-envelope wave trains discussed in Section 3.2, a high-quality fit can recover characteristic scales of the problem, like the crossing-time of the AWs through the magnetosphere ($T_{b}\approx T_{\rm bounce}$). We find that different injection scenarios – like the interacting waves analyzed in Section 3.4 – show decaying tails with less regularity and are harder to approximate by the proposed fit function. The difference in the decaying tails likely implies that late-stage dynamics are affected by the initial perturbations, especially the number of seed AW envelopes and their length.
Plastic damping of AWs during their interaction with the magnetar surface (e.g., Li & Beloborodov, 2015) can speed up the dissipation of the initially injected magnetospheric excess energy. Realistic absorption mechanisms likely remove $\gtrsim 10\%$ of AW energy in each surface interaction, where most of the AW energy is quickly absorbed and dissipated inside the star. However, the impact of crustal damping on the luminosity of outgoing FMS waves is small (see Figure 9). The produced FMS waves gain most of their energy during the first bounce of the injected AWs, as they interact around the equatorial plane. Hence, the most significant production of FMS wave energy occurs before the plastic damping acts on the reflected AWs. Electric zones of $E\approx B$ can naturally develop in parts of FMS waves leaving the inner magnetosphere with polarizations such that the wave magnetic field is directed opposite to the stellar magnetic field (Beloborodov, 2023; Chen et al., 2022b; Mahlmann et al., 2022). In the following section, we discuss the implications of these findings for FRBs.
@@ 4.1 Implications for Fast Radio Bursts
FRBs are bright millisecond-duration flashes of radio waves in the gigahertz band ($0.1-10$ GHz, CHIME/FRB Collaboration, 2019), with at least one observational association to a galactic magnetar (SGR 1935+2154 CHIME/FRB Collaboration, 2020). Between many possible generation mechanisms (reviews, e.g., in Lyubarsky, 2021; Petroff et al., 2022; Zhang, 2023), two scenarios for radio wave generation in the outer magnetar magnetosphere have been modelled consistently at relevant plasma scales: shock-mediated FRB generation at magnetized shocks by the synchrotron maser mechanism (e.g., Ghisellini, 2016; Plotnikov & Sironi, 2019; Metzger et al., 2019) or mode conversion (Thompson, 2022) on the one hand, and reconnection-mediated radio wave injection by in the compressed magnetospheric current sheet on the other hand (Lyubarsky, 2021; Mahlmann et al., 2022; Wang et al., 2023).
Reconnection-mediated FRB generation models rely on a long-wavelength FMS pulse (roughly millisecond duration) propagating to the outer magnetosphere up to the magnetar light cylinder. Mode conversion of Alfvén waves at curved field lines can generate such long pulses in two ways. First, magnetospheric crossings of long-wavelength [^4] The terms ‘short-wavelength’ and ‘long-wavelength’ can be interpreted with respect to different reference scales. For example, the wavelength $\lambda$ can be compared to the length $L$ of the field line along which a seed AW propagates. Alternatively, $\lambda$ could be compared to the scale of variation of the background magnetic field $L_{B}\sim B_{r,\rm dip}/|\text{d}B_{r,\rm dip}/\text{d}r|\sim r/3$. In Section 3.2, seed AWs have short wavelengths of $\lambda_{0}=R_{\ast}/4\lesssim L_{B}<L$. The first train of generated FMS waves has shorter wavelengths of $\lambda_{1}=R_{\ast}/8<L_{B}<L$. Later FMS modes have long wavelengths $\lambda_{2}\approx L\approx 3.5R_{\ast}>L_{B}$. These scales are separated by a factor $\lambda_{2}/\lambda_{1}\approx 30$. seed AWs generate FMS waves (Section 3.4) with conversion efficiencies $\eta$ as evaluated in Section 3.3. Second, later bounces of seed AWs result in FMS waves with modulations at scales of the Afvén-crossing-time of a tube of dipolar field lines anchored to the magnetar surface at both ends (Section 3.2). The bottom panels of Figure LABEL:fig:a1bounce lucidly illustrate this production of late-phase long-wavelength modes. The first magnetospheric crossing generates high-frequency FMS waves with wavelengths prescribed by the seed AWs. The subsequent bounce induces a long-wavelength FMS pulse with a wavelength defined by the light-crossing time of the AWs in the inner magnetosphere. For the chosen injection colatitude of $\theta_{s}=45^{\circ}$, field lines have a length of $L\approx 3.5R_{\ast}$. The light crossing time for dipolar field lines is
@@
Therefore, bounces in the inner magnetosphere ($\theta_{s}=45^{\circ}$) could provide long-wavelength FMS pulses of duration $T_{\rm bounce}\approx 0.1\,{\rm ms}$. This long-wavelength time scale can be compared to the observational properties of FRBs: their total duration ($\lesssim 1\,{\rm ms}$), their short duration sub-structures (e.g., Pastor-Marazuela et al., 2023; Snelders et al., 2023), and their quasi-periodic behavior (e.g., with hundreds of millisecond periods, see Andersen et al., 2022). AW activity in the inner magnetosphere could generate FMS waves with variability on these timee scales. Especially the pulse width would be compatible with the low-frequency FMS pulse required for reconnection-mediated FRB generation in the outer magnetosphere (see further review of feedback mechanisms by Mahlmann, 2024). Stressing field lines closer to poles will increase the bounce time. For instance, it approaches typical FRB durations of $T_{\rm bounce}\approx 1\,{\rm ms}$ as required by reconnection-mediated models for $\theta_{s}=17.5^{\circ}$. From the bottom panel of Figure LABEL:fig:a1bounce one can infer directly that for the chosen seed AWs ($b_{\rm eq}=0.2$), the injected FMS waves have amplitudes of $\delta E_{F}/B=r_{c}E^{\phi}/B\lesssim 0.05$ at $r/R_{\ast}\approx 2$. On the rapidly decaying dipole field, relative FMS wave amplitudes scale as $\delta B/B_{\rm dip}\propto r^{2}$ (e.g., Chen et al., 2022b) and develop very large amplitudes compared to the magnetic background at the magnetar light cylinder. FMS wave amplitudes as simulated in this work would surpass the compression required by reconnection-mediated FRB models by several orders of magnitude (see Mahlmann et al., 2022, Equation 28).
@@
If the FMS wave magnetic fields are directed opposite to the magnetospheric fields, FMS waves can generate electric zones with $E\approx B$. Such zones are efficient sites of plasma heating (Levinson, 2022) or particle acceleration in radiative shocks (Beloborodov, 2023; Chen et al., 2022b). For FMS waves generated close to the stellar surface, the radius at which electric zones develop for equatorial propagation is $R_{\rm nl}/R_{\ast}\approx(2b_{0})^{-1/2}$. Very low injection amplitudes $b_{0}\approx 10^{-6}$ are sufficient to create non-linear FMS waves with electric zones for $r/R_{\ast}\lesssim 10^{3}$. The wave dynamics discussed in Section 3.2 illustrate an important aspect of the electric-zone-paradigm: only those parts of the FMS waves that carry a magnetic field in the direction opposite to the background field develop $E\gtrsim B$ dynamics. Figure LABEL:fig:a1bounce shows that the corrections only act on waves with one polarity of $\delta E_{F}$. The other polarity propagates unobstructed – here carrying a high-frequency signal imprinted on the long-wavelength envelope. Besides instabilities in the inner magnetosphere (see Mahlmann et al., 2022; Mahlmann, 2024), the mode conversion mechanism discussed in this paper can be an additional source of ubiquitous (sub-GHz) FMS waves leaving the inner magnetosphere and possibly developing magnetized shocks. Then, they are viable precursors for shock-mediated FRB generation in the outer magnetosphere. FMS waves that have different polarizations would still be able to provide the compression relevant for reconnection-mediated models.
@@ 4.2 Limitations
This paper explores FFE simulations of wave propagation and dissipation in axisymmetric dipole fields to model some of the dynamics expected in highly magnetized magnetar magnetospheres. FFE simulations, though an efficient tool to model the propagation of magnetic energy in strong magnetic fields, have significant limitations. For the method employed in this work, we disclaim limitations exhaustively in Mahlmann et al. (2020a, b, 2022, 2023). Two aspects are highly relevant to the results presented above. First, our models include zones in which the electric field strength becomes comparable to the magnetic field strength ($E\approx B$). FFE methods enforce the magnetic dominance constraint $E<B$ of MHD by artificially dissipating the surplus electric field (see detailed comments by McKinney, 2006). They do not capture the plasma kinetic feedback of electromagnetic dissipation accurately. We, therefore, omit statements about the dissipation of FMS energy in electric zones but refer the reader to the relevant works in the literature (Levinson, 2022; Chen et al., 2022b; Beloborodov, 2023). Future work will be conducted with methods that include the energy transfer between electromagnetic waves and magnetospheric plasma consistently (i.e., MHD or particle-in-cell). Second, AWs propagating along curved magnetic field lines develop significant shear that can quickly reach the limits of grid diffusion (Section 3.2). In Figure 4, we evaluate the resolution requirements for resolving the FMS wave injection during a significant number of AW bounces in the inner magnetosphere. To resolve three magnetospheric crossings without significant diffusion, resolutions of $N_{r}(N_{R_{\ast}})\times N_{\theta}=19968(512)\times 1600$ are necessary, to resolve four to five bounces we need $N_{r}(N_{R_{\ast}})\times N_{\theta}=39936(1024)\times 3200$. These resolutions are very large even for axisymmetric configurations.
Constraining the magnetosphere to axisymmetry significantly limits the solution space for relativistic wave interactions (e.g., Lyubarsky, 2018; TenBarge et al., 2021; Ripperda et al., 2021; Nättilä & Beloborodov, 2022; Golbraikh & Lyubarsky, 2023). Giving accurate limits on conversion efficiencies and ultimately turbulent dissipation will require fully-fledged three-dimensional models. The provided resolution limits can guide future magnetospheric simulations targeting wave interactions. The extreme requirements for resolving many wave interactions likely make comprehensive parameter scans prohibitive. Future work should therefore include analytic theory of magnetospheric waves and localized simulations with high resolution that accurately capture the dynamics in small parts of the magnetosphere.
@@ 5 Conclusion
This work reviews key aspects of relativistic plasma wave interactions in highly magnetized magnetar magnetospheres. We conduct force-free electrodynamics (see Section 2) simulations that follow the dynamics of Afvén waves injected into the magnetosphere by shearing the magnetar surface. These boundary conditions crudely mimic the possible stresses exerted on magnetic field lines by crustal failures. AWs traverse the magnetosphere along magnetic field lines; in dipolar geometries, they bounce back and forth between the hemispheres bounded by the highly conductive stellar surface. There are several mechanisms by which AWs generate fast magnetosonic (FMS) waves that can propagate outwards, escape the inner magnetar magnetosphere, and drive energetic radiative outbursts. First, AWs propagating along curved field lines continuously inject FMS waves (Section 2.2). The coupling strength depends on the shear between adjacent field lines (i.e., the injection colatitude of seed AWs, see Figure 1) and the seed AW amplitude. Second, counter-propagating AWs can interact to drive enhanced FMS wave injection (Section 2.3). The interaction of waves with short envelopes and no phase shift shows the largest conversion efficiency: here up to three times larger than for single AWs without interactions (Section 3.3). Third, strongly sheared AWs can bounce in the inner magnetosphere. For example, long-envelope wave trains (Section 3.2) generate high-frequency FMS waves during their first crossing. Subsequently, the remaining AWs drive low-frequency FMS wave packets. Initially with large amplitudes, their strength decreases with every bounce.
The injected FMS waves could be relevant for radiative phenomena observed from magnetars. FMS waves generated in the inner magnetosphere can be ubiquitous sources of electric zones, where the strengths of electric and magnetic fields become similar. These zones could drive shocks as one promising mechanism of FRB generation. AWs propagating through the inner magnetosphere (see Figure LABEL:fig:a1bounce) could generate low-frequency magnetic perturbations and drive FRB generation in the outer magnetosphere. Understanding wave propagation and conversion in three dimensions will be key for explaining energy transport from the scale of crustal failures on the magnetar surface ($\sim 10^{5}\,{\rm cm}$) to the radiative scale (below $1\,{\rm cm}$). This work includes important steps to a more complete theory of wave dynamics in the magnetar magnetosphere.

### v2

@@ 4 Discussion
Here, $\mu_{s}\approx 10^{30}\,{\rm erg}\,{\rm cm}^{-3}$ is the shear modulus of the deep crust and $s<0.1$ denotes the elastic strain. Alfvén waves emerging from the stressed regions during the quake duration $t_{\rm Quake}$ carry a fraction $f$ of $E_{\rm Quake}$ (Bransgrove et al., 2020). Their power is:
@@
Throughout Section 3 we numerically evaluate conversion efficiencies $\eta$ for two different mode conversion mechanisms in axisymmetry: mode conversion at curved magnetic field lines and by nonlinear three-wave interactions. Typical conversion efficiencies range between $\eta\approx 1-10\%$ for low-amplitude waves; higher efficiencies can be achieved by counter-propagating interacting Alfvén waves and for large seed wave amplitudes (see Section 3.3). Directly applying these efficiencies to Equation (25) and assuming that most of the FMS wave energy is concentrated in the short initial pulse of the first magnetospheric crossing yields FMS wave luminosities of $L_{\rm FMS}\approx 10^{42}-10^{43}\,{\rm erg}/{\rm s}$. Alfvén waves can generate FMS waves of twice their frequency during the first crossing of the magnetosphere. Subsequent bounces produce more FMS waves during each crossing. These late-episode FMS waves gradually decrease in luminosity but extend the possible signal durations substantially. We describe the FMS wave tail generated by ongoing Alfvén wave activity in the inner magnetosphere by a fit function (see Appendix A). Especially for the long-envelope wave trains discussed in Section 3.2, a high-quality fit can recover characteristic scales of the problem, like the crossing time of the Alfvén waves through the magnetosphere ($T_{b}\approx T_{\rm bounce}$). We find that different injection scenarios – like the interacting waves analyzed in Section 3.4 – show decaying tails with less regularity and are harder to approximate by the proposed fit function. The difference in the decaying tails likely implies that late-stage dynamics are affected by the initial perturbations, especially the number of seed Alfvén wave envelopes and their length.
Plastic damping of Alfvén waves during their interaction with the magnetar surface (e.g., Li & Beloborodov, 2015) can speed up the dissipation of the initially injected magnetospheric excess energy. Realistic absorption mechanisms likely remove $\gtrsim 10\%$ of Alfvén wave energy in each surface interaction, where most of the Alfvén wave energy is quickly absorbed and dissipated inside the star. However, the impact of crustal damping on the luminosity of outgoing FMS waves is small (see Figure 9). The produced FMS waves gain most of their energy during the first bounce of the injected Alfvén waves, as they interact around the equatorial plane. Hence, the most significant production of FMS wave energy occurs before the plastic damping acts on the reflected Alfvén waves. Electric zones of $E\approx B$ can naturally develop in parts of FMS waves leaving the inner magnetosphere with polarizations such that the wave magnetic field is directed opposite to the stellar magnetic field (Beloborodov, 2023; Chen et al., 2022b; Mahlmann et al., 2022). In the following section, we discuss the implications of these findings for FRBs.
@@ 4.1 Implications for Fast Radio Bursts
FRBs are bright millisecond-duration flashes of radio waves in the gigahertz band ($0.1-10$ GHz, CHIME/FRB Collaboration, 2019), with at least one observational association to a galactic magnetar (SGR 1935+2154 CHIME/FRB Collaboration, 2020). Between many possible generation mechanisms (reviews, e.g., in Lyubarsky, 2021; Petroff et al., 2022; Zhang, 2023), two scenarios for radio wave generation in the outer magnetar magnetosphere have been modeled consistently at relevant plasma scales: shock-mediated FRB generation at magnetized shocks by the synchrotron maser mechanism (e.g., Ghisellini, 2016; Plotnikov & Sironi, 2019; Metzger et al., 2019) or mode conversion (Thompson, 2022) on the one hand, and reconnection-mediated radio wave injection by in the compressed magnetospheric current sheet on the other hand (Lyubarsky, 2021; Mahlmann et al., 2022; Wang et al., 2023).
Reconnection-mediated FRB generation models rely on a long-wavelength FMS pulse (roughly millisecond duration) propagating to the outer magnetosphere up to the magnetar light cylinder. Mode conversion of Alfvén waves at curved field lines can generate such long pulses in two ways. First, magnetospheric crossings of long-wavelength [^4] The terms ‘short wavelength’ and ‘long wavelength’ can be interpreted with respect to different reference scales. For example, the wavelength $\lambda$ can be compared to the length $L$ of the field line along which a seed Alfvén wave propagates. Alternatively, $\lambda$ could be compared to the scale of variation of the background magnetic field $L_{B}\sim B_{r,\rm dip}/|\text{d}B_{r,\rm dip}/\text{d}r|\sim r/3$. In Section 3.2, seed Alfvén waves have short wavelengths of $\lambda_{0}=R_{\ast}/4\lesssim L_{B}<L$. The first train of generated FMS waves has shorter wavelengths of $\lambda_{1}=R_{\ast}/8<L_{B}<L$. Later FMS modes have long wavelengths $\lambda_{2}\approx L\approx 3.5R_{\ast}>L_{B}$. These scales are separated by a factor $\lambda_{2}/\lambda_{1}\approx 30$. seed Alfvén waves generate FMS waves (Section 3.4) with conversion efficiencies $\eta$ as evaluated in Section 3.3. Second, later bounces of seed Alfvén waves result in FMS waves with modulations at scales of the Afvén-crossing time of a tube of dipolar field lines anchored to the magnetar surface at both ends (Section 3.2). The bottom panels of Figure LABEL:fig:a1bounce lucidly illustrate this production of late-phase long-wavelength modes. The first magnetospheric crossing generates high-frequency FMS waves with wavelengths prescribed by the seed Alfvén waves. The subsequent bounce induces a long-wavelength FMS pulse with a wavelength defined by the light-crossing time of the Alfvén waves in the inner magnetosphere. For the chosen injection colatitude of $\theta_{s}=45^{\circ}$, field lines have a length of $L\approx 3.5R_{\ast}$. The light-crossing time for dipolar field lines is
@@
Therefore, bounces in the inner magnetosphere ($\theta_{s}=45^{\circ}$) could provide long-wavelength FMS pulses of duration $T_{\rm bounce}\approx 0.1\,{\rm ms}$. This long-wavelength time scale can be compared to the observational properties of FRBs: their total duration ($\lesssim 1\,{\rm ms}$), their short duration substructures (e.g., Pastor-Marazuela et al., 2023; Snelders et al., 2023), and their quasi-periodic behavior (e.g., with hundreds of millisecond periods, see Andersen et al., 2022). Alfvén wave activity in the inner magnetosphere could generate FMS waves with variability on these timee scales. Especially the pulse width would be compatible with the low-frequency FMS pulse required for reconnection-mediated FRB generation in the outer magnetosphere (see further review of feedback mechanisms by Mahlmann, 2024). Stressing field lines closer to poles will increase the bounce time. For instance, it approaches typical FRB durations of $T_{\rm bounce}\approx 1\,{\rm ms}$ as required by reconnection-mediated models for $\theta_{s}=17.5^{\circ}$. From the bottom panel of Figure LABEL:fig:a1bounce one can infer directly that for the chosen seed Alfvén waves ($b_{\rm eq}=0.2$), the injected FMS waves have amplitudes of $\delta E_{F}/B=r_{c}E^{\phi}/B\lesssim 0.05$ at $r/R_{\ast}\approx 2$. On the rapidly decaying dipole field, relative FMS wave amplitudes scale as $\delta B/B_{\rm dip}\propto r^{2}$ (e.g., Chen et al., 2022b) and develop very large amplitudes compared to the magnetic background at the magnetar light cylinder. FMS wave amplitudes as simulated in this work would surpass the compression required by reconnection-mediated FRB models by several orders of magnitude (see Mahlmann et al., 2022, Equation 28).
@@
If the FMS wave magnetic fields are directed opposite to the magnetospheric fields, FMS waves can generate electric zones with $E\approx B$. Such zones are efficient sites of plasma heating (Levinson, 2022) or particle acceleration in radiative shocks (Beloborodov, 2023; Chen et al., 2022b). For FMS waves generated close to the stellar surface, the radius at which electric zones develop for equatorial propagation is $R_{\rm nl}/R_{\ast}\approx(2b_{0})^{-1/2}$. Very low injection amplitudes $b_{0}\approx 10^{-6}$ are sufficient to create nonlinear FMS waves with electric zones for $r/R_{\ast}\lesssim 10^{3}$. The wave dynamics discussed in Section 3.2 illustrate an important aspect of the electric-zone-paradigm: only those parts of the FMS waves that carry a magnetic field in the direction opposite to the background field develop $E\gtrsim B$ dynamics. Figure LABEL:fig:a1bounce shows that the corrections only act on waves with one polarity of $\delta E_{F}$. The other polarity propagates unobstructed – here carrying a high-frequency signal imprinted on the long-wavelength envelope. Besides instabilities in the inner magnetosphere (see Mahlmann et al., 2022; Mahlmann, 2024), the mode conversion mechanism discussed in this paper can be an additional source of ubiquitous (sub-GHz) FMS waves leaving the inner magnetosphere and possibly developing magnetized shocks. Then, they are viable precursors for shock-mediated FRB generation in the outer magnetosphere. FMS waves that have different polarizations would still be able to provide the compression relevant for reconnection-mediated models.
@@ 4.2 Limitations
This paper explores FFE simulations of wave propagation and dissipation in axisymmetric dipole fields to model some of the dynamics expected in highly magnetized magnetar magnetospheres. FFE simulations, though an efficient tool to model the propagation of magnetic energy in strong magnetic fields, have significant limitations. For the method employed in this work, we disclaim limitations exhaustively in Mahlmann et al. (2020a, b, 2022, 2023). Two aspects are highly relevant to the results presented above. First, our models include zones in which the electric field strength becomes comparable to the magnetic field strength ($E\approx B$). FFE methods enforce the magnetic dominance constraint $E<B$ of MHD by artificially dissipating the surplus electric field (see detailed comments by McKinney, 2006). They do not capture the plasma kinetic feedback of electromagnetic dissipation accurately. We, therefore, omit statements about the dissipation of FMS energy in electric zones but refer the reader to the relevant works in the literature (Levinson, 2022; Chen et al., 2022b; Beloborodov, 2023). Future work will be conducted with methods that include the energy transfer between electromagnetic waves and magnetospheric plasma consistently (i.e., MHD or particle-in-cell). Second, Alfvén waves propagating along curved magnetic field lines develop significant shear that can quickly reach the limits of grid diffusion (Section 3.2). In Figure 4, we evaluate the resolution requirements for resolving the FMS wave injection during a significant number of Alfvén wave bounces in the inner magnetosphere. To resolve three magnetospheric crossings without significant diffusion, resolutions of $N_{r}(N_{R_{\ast}})\times N_{\theta}=19968(512)\times 1600$ are necessary, to resolve four to five bounces we need $N_{r}(N_{R_{\ast}})\times N_{\theta}=39936(1024)\times 3200$. These resolutions are very large even for axisymmetric configurations.
Constraining the magnetosphere to axisymmetry significantly limits the solution space for relativistic wave interactions (e.g., Lyubarsky, 2018; TenBarge et al., 2021; Ripperda et al., 2021; Nättilä & Beloborodov, 2022; Golbraikh & Lyubarsky, 2023). Giving accurate limits on conversion efficiencies and ultimately turbulent dissipation will require fully fledged three-dimensional models. The provided resolution limits can guide future magnetospheric simulations targeting wave interactions. The extreme requirements for resolving many wave interactions likely make comprehensive parameter scans prohibitive. Future work should therefore include analytic theory of magnetospheric waves and localized simulations with high resolution that accurately capture the dynamics in small parts of the magnetosphere.
@@ 5 Conclusion
This work reviews key aspects of relativistic plasma wave interactions in highly magnetized magnetar magnetospheres. We conduct FFE (see Section 2) simulations that follow the dynamics of Afvén waves injected into the magnetosphere by shearing the magnetar surface. These boundary conditions crudely mimic the possible stresses exerted on magnetic field lines by crustal failures. Alfvén waves traverse the magnetosphere along magnetic field lines; in dipolar geometries, they bounce back and forth between the hemispheres bounded by the highly conductive stellar surface. There are several mechanisms by which Alfvén waves generate FMS waves that can propagate outward, escape the inner magnetar magnetosphere, and drive energetic radiative outbursts. First, Alfvén waves propagating along curved field lines continuously inject FMS waves (Section 2.2). The coupling strength depends on the shear between adjacent field lines (i.e., the injection colatitude of seed Alfvén waves, see Figure 1) and the seed Alfvén wave amplitude. Second, counter-propagating Alfvén waves can interact to drive enhanced FMS wave injection (Section 2.3). The interaction of waves with short envelopes and no phase shift shows the largest conversion efficiency: here up to three times larger than for single Alfvén waves without interactions (Section 3.3). Third, strongly sheared Alfvén waves can bounce in the inner magnetosphere. For example, long-envelope wave trains (Section 3.2) generate high-frequency FMS waves during their first crossing. Subsequently, the remaining Alfvén waves drive low-frequency FMS wave packets. Initially with large amplitudes, their strength decreases with every bounce.
The injected FMS waves could be relevant for radiative phenomena observed from magnetars. FMS waves generated in the inner magnetosphere can be ubiquitous sources of electric zones, where the strengths of electric and magnetic fields become similar. These zones could drive shocks as one promising mechanism of FRB generation. Alfvén waves propagating through the inner magnetosphere (see Figure LABEL:fig:a1bounce) could generate low-frequency magnetic perturbations and drive FRB generation in the outer magnetosphere. Understanding wave propagation and conversion in three dimensions will be key for explaining energy transport from the scale of crustal failures on the magnetar surface ($\sim 10^{5}\,{\rm cm}$) to the radiative scale (below $1\,{\rm cm}$). This work includes important steps to a more complete theory of wave dynamics in the magnetar magnetosphere.